OpenStudy (freckles):

Write 74057 as a product of two primes. No calculators please.

OpenStudy (crashonce):

split the number into its prime decomposition

OpenStudy (freckles):

How would you do that?

OpenStudy (crashonce):

try to divide it by the lowest prime

OpenStudy (crashonce):

and go up

OpenStudy (freckles):

lol that's a lot of primes

OpenStudy (mathstudent55):

It's not divisible by 2, 3, 5, 7

OpenStudy (freckles):

I will give a hint. Think quadratic expression.

OpenStudy (crashonce):

im not sure, im just trying and slooking for a pattern in the results

OpenStudy (freckles):

By the way the hint I gave is one way to do it. There might be other ways.

OpenStudy (anonymous):

quadratic expression?

OpenStudy (anonymous):

i wonder what base

OpenStudy (freckles):

it might be in terms of 10^2

OpenStudy (anonymous):

hmm

ganeshie8 (ganeshie8):

this doesn't look easy hmm \[74057 = x^2-y^2\]

OpenStudy (anonymous):

ooh ooh pick me !

OpenStudy (misty1212):

\[7x^2+40x+57=(7x+19)(x+3)\]

OpenStudy (freckles):

Girl power!

OpenStudy (misty1212):

yay for girls~

OpenStudy (misty1212):

put \(x=100\) and you find that \[74057=(700+19)(100+3)\]

ganeshie8 (ganeshie8):

thats very neat xD

OpenStudy (freckles):

I was going to say factor 72409 But I didn't like that one because it couldn't be written as a product of two primes

OpenStudy (misty1212):

then it would be a lot harder

OpenStudy (freckles):

\[72409=7 \cdot 10^4+24 \cdot 10^2+9 \\ 7 \cdot 10^4+3 \cdot 10^2+21 \cdot 10^2+9 \\ 10^2(7 \cdot 10^2+3)+3(7 \cdot 10^2+3) \\ (7 \cdot 10^2+3)(10^2+3)\] I wonder if there is a cute way to factor 703

ganeshie8 (ganeshie8):

(7x+3)(x+3)

ganeshie8 (ganeshie8):

\[703 = x^2-y^2\] \[x^2 - 703 = y^2\] since 27 is the least positive integer that makes the left hand side positive, we test x= 27,28,29,... that makes the left hand side a perfect square x = 28 gives \[28^2 - 703 = 81 = 9^2\] so x = 28 and y = 9 will work

OpenStudy (freckles):

sorry was stuffing my face that is a cute method \[703=x^2-y^2=(28-9)(28+9)=(19)(37)\]

ganeshie8 (ganeshie8):

changing the base to 9 or 8 might make life a bit easy/complicated...

OpenStudy (freckles):

sounds scary

ganeshie8 (ganeshie8):

\[(703)_{10} = (861)_9\] 7x^2+3 is not factorable but we can factor 8x^2+6x+1

OpenStudy (anonymous):

well sqrt 703 <27 :| not too much numbers to check

OpenStudy (anonymous):

and also there is a thing 703 is odd 703=2n+1=(n+1)^2-1^2 703=r 4n+3 =(2n+1)^2-1^2 or n=175 (2n+3)^2-2(n+1) failed lets try 703=8n+7--->n=87 =(8n+7)^2+(41n+16) (41n+16)=3 mod 10 failed :| 703=10n+3 --->n=7 (10n+7)^2-(1390n+6) failed 703=12n+7 >>> n=58 (12n+7)^2-(156n+42) failed 703=14n+3 (14n+3)^2-(70n+6) failed 703=16n+15-->n=43 (16n+15)^2-(464n+210) failed lol too much k since 28 works :P

OpenStudy (anonymous):

i rather to use this :- 

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

sieve of eratosthenes always works but how would you know when to give up testing ?

OpenStudy (anonymous):

when sqrt n<k such that kn+q failed